the perimeter of a rectangle is 160 ft. One fourth the length is the same as
twice the width. Find the dimensions of the rectangle.

Answer by jainenderkapoor(61)

WHAT IS THE FORMULA FOR PERIMETER OF A RECTANGLE
Perimeter = 2 (Length + Breadth)
Suppose that the width/breadth of rectangle is x ft
Then we have 1/4 Length = 2x
Length = 2x * 1/4
Length = x/2
Now Perimeter = 2 (Length + Breadth) = 160
2 (x/2 + x) = 160
2 (3x/2) = 160
3x = 160
x = 160/3
So Width/Breadth = 160/3 ft
Length = (160/3)/2 = 160/6
CHECK = Perimeter = 2 (Length + Breadth)
= 2 (160/6 + 160/3)
= 2 *160(1/6 + 1/3)
= 2 * 160((2+1)/6)
= 2 * 160 * 3 /6
= 160 ft

HOPE THIS IS CLEAR TO YOU. IN CASE YOU HAVE ANY DOUBT PLEASE CONTACT ME.
I AM AVAILBLE FOR ONLINE TUTORING ALSO. MY EMAIL IS kapoorjai1@rediffmail.com